Cylinder subjected to internal pressure with an internal diameter of 350 mm and a 25 mm wall thickness. For the cylinder material, KIC = 80 MPa√m, Sy = 1172 MPa, and Sut = 1324 MPa. If the cylinder contains a radial crack in the longitudinal direction of depth 12.5, mm determine the pressure that will cause uncontrollable crack growth.


1
Expert's answer
2022-02-06T12:31:23-0500

Pressure that will cause uncontrollable crack growth.

"P =\frac{K_{IC}\times A}{\beta\times \sqrt{\pi a}} =\frac{80\times 14.23\times 10^3 }{1.3\times \sqrt{\pi 25}} =98.81" kN